# Security Concerns for Large Language Models: A Survey

### v2

@@ VI-C Limitations and Insufficiencies of Current Defenses
Critically, the findings by Hubinger et al.[12] suggest that current safety training paradigms, including adversarial training and RLHF, may be insufficient to remove deeply embedded deceptive behaviors or “sleeper agent” capabilities. If models can learn to strategically appear aligned during training and evaluation, they may create a dangerous false impression of safety. This underscores the argument that ensuring an agentic AI will not cause harm is a notoriously difficult, unsolved technical problem, especially as current training methods for agentic AI can lead to issues such as goal misspecification and misgeneralization[35]. As one survey notes, most defense mechanisms “present inherent limitations that can be exploited by informed and sufficiently skilled malicious actors”[11]. In summary, while a multi-layered approach (sanitization, monitoring, aligned training, human oversight) can mitigate risk, no single defense is fully effective against the evolving threat landscape, particularly the challenge of ensuring genuine alignment and preventing strategic deception in advanced agents[11, 12, 13], leading some researchers to propose fundamentally different, non-agentic AI paradigms such as “Scientist AI”, designed for trustworthiness and safety from the ground up by focusing on understanding and probabilistic inference rather than goal pursuit [35]. However, the “Scientist AI” paradigm remains in its infancy, and it is still unclear whether such non-agentic systems can ultimately match the general intelligence demonstrated by agentic AI models.
TABLE V: Overview of Defense Mechanisms Against LLM Threats and Their Limitations
@@
Detection Runtime Oversight External layers monitoring and intervening in agent plans. Harmful actions by agents. Defining ”harmful” is hard; complex implementation. Wang et al. [47], Crouse et al. [48]
VII Open Challenges and Future Directions
@@
VII-A Adaptive and Automated Attacks
@@
VII-B Robust Alignment, Verification, and Control of Agentic LLMs
This is perhaps the most critical and challenging area. How can we guarantee that an LLM truly understands, internalizes, and adheres to complex human intentions, especially when it possesses advanced reasoning and planning capabilities? Current alignment techniques (e.g., RLHF, adversarial training) have shown limitations against strategic deception and “sleeper agents” [12, 13]. New paradigms are needed for:
@@ VII-B Robust Alignment, Verification, and Control of Agentic LLMs
VII-C Data Integrity and Provenance
@@
VII-D Detection of Malicious Uses and Content
@@ VII-E Standardization and Collaboration
VII-F Human-AI Interaction Research
@@
VII-G Emerging Proactive Measures and Ongoing Efforts

### v3

@@ 6.3 Limitations and Insufficiencies of Current Defenses
Critically, the findings by Hubinger et al.[21] suggest that current safety training paradigms, including adversarial training and RLHF, may be insufficient to remove deeply embedded deceptive behaviors or “sleeper agent" capabilities. If models can learn to strategically appear aligned during training and evaluation, they may create a dangerous false impression of safety. This underscores the argument that ensuring an agentic AI will not cause harm is a notoriously difficult, unsolved technical problem, especially as current training methods for agentic AI can lead to issues such as goal misspecification and misgeneralization[6]. As one survey notes, most defense mechanisms “present inherent limitations that can be exploited by informed and sufficiently skilled malicious actors”[14]. In summary, while a multi-layered approach (sanitization, monitoring, aligned training, human oversight) can mitigate risk, no single defense is fully effective against the evolving threat landscape, particularly the challenge of ensuring genuine alignment and preventing strategic deception in advanced agents[14, 21, 29], leading some researchers to propose fundamentally different, non-agentic AI paradigms such as “Scientist AI", designed for trustworthiness and safety from the ground up by focusing on understanding and probabilistic inference rather than goal pursuit [6]. However, the “Scientist AI” paradigm remains in its infancy, and it is still unclear whether such non-agentic systems can ultimately match the general intelligence demonstrated by agentic AI models.
Table 4: Overview of Defense Mechanisms Against LLM Threats and Their Limitations
@@
Prevention Deliberative Alignment A training paradigm that teaches a model to explicitly recall and reason over safety specifications via CoT before answering. Jailbreaks, harmful content generation, over-refusals, poor out-of-distribution generalization. Relies on the quality of the safety specifications and the judge model; effectiveness against strategic deception (scheming) is an open question. Guan et al. [18]
Prevention Instruction Hierarchy Training LLMs to prioritize instructions by source (system > user > tool) to ignore lower-privilege malicious commands. Prompt injection (direct & indirect), system prompt extraction, some jailbreaks. Can lead to "over-refusals" (ignoring benign instructions); may be too rigid if not carefully tuned. Wallace et al. [41]
@@
Detection Runtime Oversight External layers monitoring and intervening in agent plans. Harmful actions by agents. Defining "harmful" is hard; complex implementation. Wang et al. [43], Crouse et al. [11]
7 Open Challenges and Future Directions
@@
7.1 Adaptive and Automated Attacks
@@
7.2 Robust Alignment, Verification, and Control of Agentic LLMs
This is perhaps the most critical and challenging area. How can we guarantee that an LLM truly understands, internalizes, and adheres to complex human intentions, especially when it possesses advanced reasoning and planning capabilities? Current alignment techniques (e.g., RLHF, adversarial training) have shown limitations against strategic deception and “sleeper agents" [21, 29]. New paradigms are needed for:
@@ 7.2 Robust Alignment, Verification, and Control of Agentic LLMs
7.3 Data Integrity and Provenance
@@
7.4 Detection of Malicious Uses and Content
@@ 7.5 Standardization and Collaboration
7.6 Human-AI Interaction Research
@@
7.7 Emerging Proactive Measures and Ongoing Efforts
@@ 8 Conclusion
Acknowledgment
The research is supported in part by NSERC Discovery Grants (RGPIN-2024-04087), NSERC Collaborative Research and Training Experience (CREATE-554764-2021), and Canada Research Chairs Program (CRC-2019-00041).